OpenStudy (anonymous):

I don't understand how to do this problem, could someone please explain it to me? 1) Solve for x. Round to the nearest tenth. cos(62 degrees)=4/x

OpenStudy (mathstudent55):

cos 62 is simply a number.

OpenStudy (doc.brown):

a=b/x xa=xb/x xa=b xa/a=b/a x=b/a or in your case x=4/cos(62)

OpenStudy (anonymous):

So is x=4/cos(62) the answer?

OpenStudy (doc.brown):

do you have a calculator?

OpenStudy (anonymous):

Yes.

OpenStudy (doc.brown):

find the cos button

OpenStudy (anonymous):

you multiply x by both sides it candles out on the right side then you divide by cos62 so you have x=4/cos62 just plug 4/cos 62 into you calculator

OpenStudy (anonymous):

Found it.

OpenStudy (anonymous):

Okay, I understand. Thank you both for the help!!

OpenStudy (mathstudent55):